The uniform cross section of a cabin storage compartment on a jet plane is a sector of a circle with radius 55cm, and angle Ө c .  The container is 96 cm wide. The volume of the container is 0.2178 m3. What is the angle Өc




1
Expert's answer
2021-05-07T12:07:35-0400

We have given that,


"r= 55cm = 0.55m"


Length of wide "= 96cm = 0.96m"


Volume = "0.2178 m^3"


"\theta_c" = ?

We know volume of the sector

"= \dfrac{\theta_c \times \pi r^2}{360} \times length"


"0.2178 = \dfrac{\theta_c\times 3.14 \times 0.55 \times 0.55}{360} \times 0.96"


"\theta_c = \dfrac{78.408}{0.911} = 86.06"